4x-12y=6
15x4y=4
and
2x+5y=33.5
3x-10y=38

Answer by mananth(16949)

2 x + 5 y = 33.3 .............1
3 x -10 y = 38 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
4 x 10 y = 66.6
3 x -10 y = 38
Add the two equations
7 x = 104.6
/ 7
x = 14.94
plug value of x in (1)
2 x + 5 y = 33.3
29.89 + 5 y = 33.3
5 y = 33.3 -29.89
5 y = 3.41
y = 0.68

first system typo